### v2

# bibliography.bib

Abstract.
[…]
The authors thank Galen Dorpalen-Barry, Theo Douvropoulos, and Raman Sanyal for helpful discussions.
1. Introduction and main results
1.1. Background
Let $\mathcal{A}$ be a central and essential hyperplane arrangement in $V\cong\mathbb{R}^{n}$. Its lattice of flats $\mathcal{L}$ is given by the intersections of hyperplanes in $\mathcal{A}$ ordered by reverse inclusion, so that its unique bottom element is $V$ and its unique top element is $\mathbf{0}$. The chain Poincaré series of a chain $\mathcal{C}=\{C_{1}<C_{2}<\dots<C_{k}\}$ in $\mathcal{L}$ is given by the product of the Poincaré series of the intervals, with bottom and top attached,
$\operatorname{Poin}_{\mathcal{C}}(y)=\prod_{i=0}^{k}\operatorname{Poin}([C_{i},C_{i+1}];y)\,,$
where we set $C_{0}=V$ and $C_{k+1}=\mathbf{0}$. The coarse flag Hilbert-Poincaré series is defined by Maglione–Voll in [MaglioneVoll] as
$\operatorname{cfHP}_{\mathcal{A}}(y,t)=\sum_{\mathcal{C}}\operatorname{Poin}_{\mathcal{C}}(y)\cdot\left(\frac{t}{1-t}\right)^{\#\mathcal{C}}=\frac{\mathcal{N}_{\mathcal{A}}(y,t)}{(1-t)^{n}}\,,$
where the sum ranges over all chains in the intersection poset $\mathcal{L}$ not containing the bottom element $V$. Its numerator polynomial has nonnegative coefficients [poincareextended] and the Chow polynomial of the associated matroid is given by the evaluation $\operatorname{cfHP}_{\mathcal{A}}(-x,x)$ [stump2024chowaugmentedchowpolynomials]. It is moreover shown in [KühneMaglione] for simplicial arrangements that
$\mathcal{N}_{\mathcal{A}}(1,t)=\#\mathcal{R}\cdot\operatorname{E}_{n}(t)\,,$
where $\operatorname{E}_{n}(t)=\sum_{\sigma\in\operatorname{Sym}(n)}t^{\operatorname{des}(\sigma)}$ is the $n$-th Eulerian polynomial, and where $\mathcal{R}$ denotes the set of regions of the arrangement $\mathcal{A}$ given by the connected components of its complement. This raises the following natural question.
Question 1.1.
Find a companion statistic $\operatorname{st}:\mathcal{R}\times\operatorname{Sym}(n)\rightarrow\mathbb{N}$ for the descent number on $\operatorname{Sym}(n)$ such that
$\mathcal{N}_{\mathcal{A}}(y,t)=\sum_{\begin{subarray}{c}R\in\mathcal{R}\\ \sigma\in\operatorname{Sym}(n)\end{subarray}}y^{\operatorname{st}(R,\sigma)}\cdot t^{\operatorname{des}(\sigma)}\,.$
If $\mathcal{A}$ is the reflection arrangement of a real reflection group $W$, the regions are in bijection with the elements of the group $W$, so this becomes the question of finding a companion statistic on ${W\times\operatorname{Sym}(n)}$. We emphasize that it already appears to be a challenging task to find such a companion statistic, and that this statistic would in particular provide a new approach towards generalizing left-to-right minima on permutations to real reflection groups, see Remark 1.5.
The purpose of this paper is to introduce a new statistic on permutations which answers 1.1 for the symmetric group.
Let $\mathcal{B}_{n}$ be the braid arrangement given by the hyperplanes $H_{ij}=\{x_{i}=x_{j}\}$ for $11\leq i<j\leq n+$ 1 in $V=\big{\{}(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\mid x_{1}+\dots+x_{n+1}=0\big{\}}\cong\mathbb{R}^{n}$, the corresponding reflection group is $W={\mathfrak{S}}_{n+1}$. Its lattice of flats, denoted by $\Pi_{n}$, is given by the lattice of set partitions of $\{1,\dots,n+1\}$ of rank $n$, with cover relations being given by merging two blocks. Its minimal element is the all-singletons set partition and its maximal element is the all-in-one-block set partition. A set composition is a set partition together with a linear ordering of the blocks. We write $B_{1}|B_{2}|\cdots|B_{k}$ for the set composition with the blocks being ordered from left to right. The face poset of the braid arrangement $\mathcal{B}_{n}$, denoted by $\Sigma_{n}$, is given by the poset of set compositions of $\{1,\dots,n+1\}$ of rank $n$, with cover relations being given by merging two consecutive blocks. Its minimal elements are thus in natural bijection with permutations, and its unique maximal element is the all-in-one-block set composition.
Example 1.2.
The braid arrangement $\mathcal{B}_{2}$ is given below with its regions labeled by $w\in{\mathfrak{S}}_{3}$ in one-line-notation.
[…]
The element $w\in{\mathfrak{S}}_{3}$ encodes the region $R_{w}=\left\{x\in V\;|\;x_{w(1)}\leq x_{w(2)}\leq x_{w(3)}\right\}$. Its lattice of flats $\mathcal{L}$ is the lattice of set partitions of the set $\{1,2,3\}$. Hereby, integers $i,j$ are in the same block of the corresponding partition if and only if $x_{i}=x_{j}$.
$\boldsymbol{0}$ $H_{13}$ $H_{12}$ $H_{23}$cong $\bigl{\{}\{1,2\},\{3\}\bigr{\}}$$\bigl{\{}\{1,3\},\{2\}\bigr{\}}$$\bigl{\{}\{2,3\},\{1\}\bigr{\}}$$\bigl{\{}1,2,3\bigr{\}}$ $V$$\bigl{\{}\{1\},\{2\},\{3\}\bigr{\}}$
The chains in $\mathcal{L}\setminus\{V\}$ are given by $\emptyset$, $\{\boldsymbol{0}\}$, $\{H_{ij}\}$, and $\{H_{ij}<\boldsymbol{0}\}$ for $1\leq i<j\leq 3$. We obtain
$\mathcal{C}$ $\operatorname{Poin}_{\mathcal{C}}(y)$$\#\mathcal{C}$
$\emptyset$ $1+3y+2y^{2}$ 0
$\{\boldsymbol{0}\}$ $1+3y+2y^{2}$ 1
$\{H_{ij}\}$ $(1+y)^{2}$ 1
$\{H_{ij}<\boldsymbol{0}\}$ $(1+y)^{2}$ 2
$\operatorname{cfHP}_{\mathcal{B}_{2}}(y,t)=$ $\frac{(1+3y+2y^{2})+(2+3y+y^{2})t}{(1-t)^{2}}\,.$
The numerator polynomial $\mathcal{N}_{\mathcal{B}_{2}}(y,t)=(1+3y+2y^{2})+(2+3y+y^{2})t$ evaluated in $y=1$ is
$\mathcal{N}_{\mathcal{B}_{2}}(1,t)=6+6t=6\cdot(1+t)=6\cdot\operatorname{E}_{2}(t)\,.$
1.2. Main result
Let $(w,\sigma)\in{\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)$ and write $w=[\ w_{1}\ |\ w_{2}\ |\ \dots\ |\ w_{n+1}\ ]$ in one-line notation with bars labeled from left to right by the numbers $1$ through $n$ between the entries. For the given pair $(w,\sigma)$, define a maximal chain in $\Sigma_{n}$ by starting with the minimal element given by the all-singleton composition $w$, and by merging blocks by deleting the bars in the positions in the order given by $\sigma(1),\sigma(2),\dots,\sigma(n)$. For $i\in\{1,\dots,n\}$, one thus has the cover relation in $\Sigma_{n}$ given by merging blocks $B|B^{\prime}$ into $B\cup B^{\prime}$ at the bar in position $\sigma(i)$. The statistic
$\operatorname{ino}:{\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)\rightarrow\mathbb{N}$
is then defined to be the number of such indices $i$ for which the minimal elements in $B$ and in $B^{\prime}$ come in order, meaning that $\min(B)<\min(B^{\prime})$.
Example 1.3.
The poset of set compositions of the set $\{1,2,3\}$ is given below. Its minimal elements are ${\mathfrak{S}}_{3}$ and its maximal chains are in bijection with ${\mathfrak{S}}_{3}\times\operatorname{Sym}(2)$. The chain corresponding to $w=123$ and $\sigma=21$ is marked by nonviolent violet thick edges and the chain corresponding to $w=312$ and $\sigma=12$ is marked by golden ratio orange dashed edges.
$123$ $3|12$ $13|2$ $1|23$ $12|3$ $23|1$ $2|13$ $1|2|3$ $1|3|2$ $3|1|2$ $3|2|1$ $2|3|1$ $2|1|3$
The merging blocks $3|1$ in the cover relation $3|1|2\prec 13|2$ are not in order, while the merging blocks in $13|2\prec 123$ are. Thus, $\operatorname{ino}({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})=1$. All the blocks in the chain $({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})$ come in order, so $\operatorname{ino}({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})=2$.
Theorem 1.4.
For the braid arrangement $\mathcal{B}_{n}$, we have
$\mathcal{N}_{\mathcal{B}_{n}}(y,t)=\sum_{\begin{subarray}{c}w\in{\mathfrak{S}}_{n+1}\\ \sigma\in\operatorname{Sym}(n)\end{subarray}}y^{\operatorname{ino}(w,\sigma)}\cdot t^{\operatorname{des}(\sigma)}\,.$
We further refine this theorem in the remainder of this introduction and then prove it in Section 2. Beside its natural appearance, the motivation for studying this companion statistic for the symmetric group is its connection to left-to-right minima for general real reflection groups.
Remark 1.5 (Motivation).
For a real reflection group $W$, 1.1 becomes the question of providing a companion statistic for descents on $\operatorname{Sym}(n)$ in terms of the group $W$,
$\mathcal{N}_{\mathcal{A}}(y,t)=\sum_{\begin{subarray}{c}w\in W\\ \sigma\in\operatorname{Sym}(n)\end{subarray}}y^{\operatorname{st}(w,\sigma)}\cdot t^{\operatorname{des}(\sigma)}\,.$
It is known that $\mathcal{N}_{\mathcal{A}}(y,0)=\operatorname{Poin}(\mathcal{A},y)$, where the latter is the Poincaré polynomial of the arrangement, see [MaglioneVoll]. For reflection arrangements $\mathcal{A}_{W}$, this corresponds to
$\operatorname{Poin}(\mathcal{A}_{W},y)=\sum_{w\in W}y^{\ell_{R}(w)}\,,$
where $\ell_{R}:W\rightarrow\mathbb{N}$ denotes the reflection length on $W$ given by the minimal number $k$ such that $w=r_{1}\cdots r_{k}$ for reflections $r_{1},\dots,r_{k}\in\leavevmode\nobreak\ W$. We thus obtain that the yet-to-be-found companion statistic on $W\times\operatorname{Sym}(n)$ needs to be equidistributed with the reflection length on the slice $W\times\{\operatorname{id}\}$ for the identity permutation $\operatorname{id}=[1,\dots,n]\in\operatorname{Sym}(n)$,
$\sum_{w\in W}y^{\operatorname{st}(w,\operatorname{id})}=\sum_{w\in W}y^{\ell_{R}(w)}\,.$
For the symmetric group, it is immediate from the definition that the companion statistic $\operatorname{ino}(w,\operatorname{id})$ records exactly the non-left-to-right minima of the permutation $w$. It is therefore a generalization of non-left-to-right minima. For general $W$, a companion statistic $\operatorname{st}:W\times\operatorname{Sym}(n)$ would thus necessarily be a version of generalized non-left-to-right minima that is equidistributed on $W\times\{\operatorname{id}\}$ with reflection length.
1.3. Refining the main result
In order to prove Theorem 1.4, we make use of a formula for the numerator polynomial $\mathcal{N}_{\mathcal{A}}(y,t)$ based on an $R$-labeling (see Section 2.1) on the lattice of flats of $\mathcal{B}_{n}$. The set partition lattice $\Pi_{n}$ has an $R$-labeling $\lambda$ given by labeling a cover relation merging two blocks $B$ and $B^{\prime}$ by ````, see [MR0570784, Example 2.9]. We extend this $R$-labeling from the set partition lattice to an edge-labeling of the poset of set compositions $\Sigma_{n}$ by defining $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$ for $w\in{\mathfrak{S}}_{n+1}$ and $\sigma\in\operatorname{Sym}(n)$ by
(1.1) $\lambda_{i}=\varepsilon(B|B^{\prime})\cdot\max\big{\{}\min(B),\min(B^{\prime})\big{\}}$
where $B,B^{\prime}$ are the merging blocks at the $i$-th edge of the chain $(w,\sigma)$, and $\varepsilon(B|B^{\prime})\in\{\pm 1\}$ is positive if $\min(B)<\min(B^{\prime})$ and negative otherwise. In particular,
(1.2) $\operatorname{ino}(w,\sigma)=\Big{|}\big{\{}\lambda_{i}\in\lambda(w,\sigma)\mid\lambda_{i}>0\big{\}}\Big{|}\,.$
Since these two edge-labelings depend only on the merging blocks, the natural surjection from $\Sigma_{n}$ to $\Pi_{n}$ (given by forgetting the order of the blocks) maps an edge label to its absolute value. If not explicitly stated otherwise, we denote by $\lambda$ the signed version of the edge-labeling. We obtain the following proposition that we prove in Section 2.1.
Proposition 1.6.
For the braid arrangement $\mathcal{B}_{n}$, we have
$\mathcal{N}_{\mathcal{B}_{n}}(y,t)=\sum_{\begin{subarray}{c}w\in{\mathfrak{S}}_{n+1}\\ \sigma\in\operatorname{Sym}(n)\end{subarray}}y^{\operatorname{ino}(w,\sigma)}\cdot t^{\operatorname{asc}(\lambda(w,\sigma))}\,,$
where $\operatorname{asc}\big{(}\lambda(w,\sigma)\big{)}=\big{|}\{i\mid\lambda_{i}<\lambda_{i+1}\}\big{|}$ is the number of ascents of $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$.
This description looks similar to Theorem 1.4. But this is not quite the case since here, both statistics depend on $w$ and $\sigma$, so this is not a companion statistic for descents (or ascents) on $\operatorname{Sym}(n)$. We in particular emphasize that we do not have a bijection $\varphi$ on ${\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)$ sending $\operatorname{asc}\big{(}\lambda(w,\sigma)\big{)}$ to $\operatorname{des}\big{(}\varphi(\sigma)\big{)}$ while leaving $\operatorname{ino}(w,\sigma)=\operatorname{ino}\big{(}w,\varphi(\sigma)\big{)}$ invariant.
Example 1.7.
For $\Sigma_{2}$, we obtain the edge-labeling
[…]
(Here and below, we often write $\overline{x}=-x$ for $x\in\mathbb{Z}$.) The sequences $\lambda(w,\sigma)$ for the chains $({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})$ and $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})$ are in particular
$\lambda({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})=(3,2)\quad\text{and}\quad\lambda({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}312},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})=(\overline{3},2)\,.$
We obtain $\operatorname{asc}\big{(}\lambda({\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}123},{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}21})\big{)}=0$ and $\operatorname{asc}\big{(}\lambda({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}231},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}12})\big{)}=1$.
The following theorem refines Theorem 1.4.
Theorem 1.8.
Let $w\in{\mathfrak{S}}_{n+1}$, then
$\sum_{\sigma\in\operatorname{Sym}(n)}y^{\operatorname{ino}(w,\sigma)}\cdot t^{\operatorname{asc}(\lambda(w,\sigma))}$ $=\sum_{\sigma\in\operatorname{Sym}(n)}y^{\operatorname{ino}(w,\sigma)}\cdot t^{\operatorname{des}(\sigma)},$
We prove a multivariate version of this result as stated in Theorem 2.2 in Section 2.
Remark 1.9.
The right-hand side of Theorem 1.4 depends on multiple choices. The “max-of-min” labeling in (1.1) could have also been “min-of-max”, the “in-order” statistic $\operatorname{ino}(w,\sigma)$ could have been an “out-of-order” statistic, and descents on $\sigma$ could have been ascents. While multiple combinations yield the same generating function in Theorem 1.4, the given version is the unique choice for which the refinement in Theorem 1.8 holds.
Remark 1.10.
As written in [MR0570784, Example 2.9], the “max-of-min” labeling in (1.1) is the edge-labeling given in [MR0354473, Proposition 2.4] for supersolvable lattices, using the $\mathcal{M}$-chain
[…]
in $\Pi_{n}$. This $\mathcal{M}$-chain induces the admissible map $H_{ij}\mapsto j$, which in turn induces the “max-of-min” $R$-labeling. This admissible map, and thus the “max-of-min” $R$-labeling, may as well be obtained by considering the induced edge-labeling of an admissible map given in [MR0354473, Proposition 2.2] for geometric lattices, using the total order
$H_{12}<H_{13}<H_{23}<H_{14}<H_{24}<H_{34}<\dots<H_{1(n+1)}<\dots<H_{n(n+1)}$
of the atoms in $\Pi_{n}$. It is easy to check that no two labels $H_{ij}$ and $H_{i^{\prime}j}$ can occur along a maximal chain in $\Pi_{n}$, implying again that $H_{ij}\mapsto j$ is an admissible map.
2. Proofs
In this section, we
  1. 1.
  2. 2.
    provide the most refined version of the main result in Theorem 2.2 (Section 2.2),
  3. 3.
    collect several properties of the $R$-labeling and reestablish a symmetry of $\mathcal{N}_{\mathcal{B}_{n}}$ (Section 2.3),
  4. 4.
    define a poset associated to $w\in{\mathfrak{S}}_{n+1}$ and certain subsets $Y\subseteq\{1,\dots,n\}$ together with an injective vertex-labeling (Section 2.4),
  5. 5.
    use reverse $(P,\omega)$-partitions to prove the main result (Section 2.5).
2.1. Rewriting the numerator polynomial
Let $P$ be a finite and ranked poset of rank $n$ with unique minimal element $\widehat{0}$ and unique maximal element $\widehat{1}$. Denote its set of cover relations by $E(P)=\left\{(x,y)\;|\;x\prec y\right\}$. An edge-labeling $\lambda:E(P)\to\mathbb{N}_{+}$ is an $\boldsymbol{R}$-labeling if every interval in $P$ contains a unique $\lambda$-weakly increasing maximal chain. Throughout this section, we assume the edge-labeling to be injective along maximal chains. For a maximal chain =$\mathcal{M}=\{\mathcal{M}_{0}\prec\dots\prec\mathcal{M}_{n}\}$, define its labeling as $\lambda(\mathcal{M})=\big{(}\lambda(\mathcal{M}_{0},\mathcal{M}_{1}),\dots,\lambda(\mathcal{M}_{n-1},\mathcal{M}_{n})\big{)}$. For a subset $Y\subseteq\{1,\dots,n\}$, we slightly modify $\lambda(\mathcal{M})$ to obtain $\lambda(\mathcal{M},Y)=\big{(}\!\pm\lambda(\mathcal{M}_{0},\mathcal{M}_{1}),\dots,\pm\lambda(\mathcal{M}_{n-1},\mathcal{M}_{n})\big{)}$ where the sign is positive if $i\in Y$ and negative if $i\notin Y$.
The lattice of flats $\mathcal{L}$ of a hyperplane arrangement $\mathcal{A}$ always admits an $R$-labeling $\lambda$ [MR2868112, Example 3.14.5]. The numerator polynomial $\mathcal{N}_{\mathcal{A}}$ was shown in [poincareextended] to have the description
(2.1) $\mathcal{N}_{\mathcal{A}}(y,t)=\sum_{\mathcal{M},Y}y^{\#Y}\cdot t^{\operatorname{asc}(\lambda(\mathcal{M},Y))}$
where the sum ranges over all maximal chain $\mathcal{M}$ in $\mathcal{L}$ and all subsets $Y\subseteq\{1,\dots,n\}$ for $n=\operatorname{rk}(\mathcal{L})$, and where $\operatorname{asc}\big{(}\lambda(\mathcal{M},Y)\big{)}$ denotes the number of ascents of $\lambda(\mathcal{M},Y)$. We refer to [poincareextended] for further details.
Example 2.1.
Label the edges in the lattice of flats of the braid arrangement $\mathcal{B}_{2}$ by the $R$-labeling $\lambda$ defined above. The table on the right gives the relation between the signed labels along the three maximal chains $\mathcal{M}_{ij}=\{V\prec H_{ij}\prec\boldsymbol{0}\}$, depending on the subsets $Y\subseteq\{1,2\}$.
$\boldsymbol{0}$ $H_{13}$ $H_{12}$ $H_{23}$ $V$ 333322222222222233333333
$Y$emptyset $\{1\}$$\{2\}$$\{1,2\}$
$\mathcal{M}_{12}$ $-2>-3$ $2>-3$ $-2<3$ $2<3$
$\mathcal{M}_{13}$ $-3<-2$ $3>-2$ $-3<2$ $3>2$
$\mathcal{M}_{23}$
Using (2.1), the numerator polynomial for $\mathcal{B}_{2}$ is
$\mathcal{N}_{\mathcal{B}_{2}}(y,t)=y^{0}\cdot(t+2)+y^{1}\cdot(3t+3)+y^{2}\cdot(2t+1)=(1+3y+2y^{2})+(2+3y+y^{2})t$
in agreement with Example 1.2.
For a simplicial arrangement $\mathcal{A}$, the set ×$\mathcal{R}\times\operatorname{Sym}(n)$ is in natural bijection with maximal chains in the face poset $\Sigma$ of $\mathcal{A}$. The elements of this poset are the faces of $\mathcal{A}$, i.e., the cones in the simplicial fan defined by $\mathcal{A}$, and the order is reverse inclusion of cones. The minimal elements of $\Sigma$ are in particular the closures of the regions $\mathcal{R}$. Sending such a maximal chain in the face poset to its corresponding maximal chain in the lattice of flats gives a natural bijection
(2.2) ×$\mathcal{R}\times\operatorname{Sym}(n)\ \tilde{\longleftrightarrow}\ \big{\{}\text{maximal chains in }\mathcal{L}\big{\}}\times\big{\{}\text{subsets of }\{1,\dots,n\}\big{\}}\,.$
We have already seen this correspondence in the case of the braid arrangement $\mathcal{B}_{n}$. The bijection between ${\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)$ and maximal chains in $\Sigma_{n}$ was described in the paragraph before and inside Example 1.3. Moreover, sending a set composition to its corresponding set partition yields the subset $Y\subseteq\{1,\dots,n\}$ given by $Y=\big{\{}i\in\{1,\dots,n\}\mid\lambda_{i}\text{ positive}\big{\}}$ for $\lambda_{i}$ as given in (1.1).
Proof of Proposition 1.6.
Consider the bijection in (2.2) in the case of the braid arrangement as described in the previous paragraph. Let now $(w,\sigma)\in{\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)$, and let $(\mathcal{M},Y)$ be its image under the bijection. Then, by construction, we have $\operatorname{ino}(w,\sigma)=\#Y$ and at the same time $\lambda(w,\sigma)=\lambda(\mathcal{M},Y)$. ∎
2.2. Further refinement of the main result
For $(w,\sigma)\in{\mathfrak{S}}_{n+1}\times\operatorname{Sym}(n)$, let
$\operatorname{Ino}(w,\sigma)=\left\{\lambda_{i}\in\lambda(w,\sigma)\;|\;\lambda_{i}>0\right\}$
denote the set of positive integers in $\lambda(w,\sigma)$. We have seen this set already in (1.2) where we observe that $\operatorname{ino}(w,\sigma)=|\operatorname{Ino}(w,\sigma)|$. Moreover, let $\operatorname{Des}(\sigma)=\left\{i\in\{1,\dots,n-1\}\;\big{|}\;\sigma(i)>\sigma(i+1)\right\}$ denote the descent set of $\sigma$ and similarly, let $\operatorname{Asc}(\lambda\big{(}w,\sigma)\big{)}$ denote the ascent set of $\lambda(w,\sigma)$. We finally consider refined indeterminates $y_{1},\dots,y_{n},t_{1},\dots,t_{n-1}$ and write
$\boldsymbol{y}^{Y}=\prod_{i\in Y}y_{i}\,,\quad\boldsymbol{t}^{T}=\prod_{i\in T}t_{i}$
for subsets $Y\subseteq\{1,\dots,n\}$ and $T\subseteq\{1,\dots,n-1\}$.
Theorem 2.2.
Let $w\in{\mathfrak{S}}_{n+1}$, then
$\sum_{\sigma\in\operatorname{Sym}(n)}\boldsymbol{y}^{\operatorname{Ino}(w,\sigma)}\cdot\boldsymbol{t}^{\operatorname{Asc}(\lambda(w,\sigma))}=\sum_{\sigma\in\operatorname{Sym}(n)}\boldsymbol{y}^{\operatorname{Ino}(w,\sigma)}\cdot\boldsymbol{t}^{\operatorname{Des}(\sigma)}\,.$
Example 2.3.
In our running example $\mathcal{B}_{2}$, we have
$w$ $\lambda(w,12)$ $\lambda(w,21)$ $\boldsymbol{y}^{\operatorname{Ino}(w,12)}$ $\boldsymbol{y}^{\operatorname{Ino}(w,21)}$ $\boldsymbol{t}^{\operatorname{Asc}(\lambda(w,12))}$ $\boldsymbol{t}^{\operatorname{Asc}(\lambda(w,21))}$ $\boldsymbol{t}^{\operatorname{Des}(12)}$ $\boldsymbol{t}^{\operatorname{Des}(21)}$
123 23 32 $y_{2}y_{3}$ $y_{2}y_{3}$ $t_{1}$ $1$ $1$ $t_{1}$
132 32 $\overline{3}2$ $y_{2}y_{3}$ $y_{2}$ $1$ $t_{1}$ $1$ $t_{1}$
213 $\overline{2}3$ $3\overline{2}$ $y_{3}$ $y_{3}$ $t_{1}$ $1$ $1$ $t_{1}$
231 $3\overline{2}$ $\overline{3}\overline{2}$ $y_{3}$ $1$ $1$ $t_{1}$ $1$ $t_{1}$
312 $\overline{3}2$ $2\overline{3}$ $y_{2}$ $y_{2}$ $t_{1}$ $1$ $1$ $t_{1}$
321 $\overline{3}\overline{2}$ $\overline{2}\overline{3}$ $1$ $1$ $t_{1}$ $1$ $1$ $t_{1}$
To give one concrete entry, the equation in Theorem 2.2 for $w=213$ is
$y_{3}t_{1}+y_{3}=y_{3}+y_{3}t_{1}\,.$
2.3. Behavior of the edge-labeling $\boldsymbol{\lambda}$
Throughout this section, fix $w=w_{1}\dots w_{n+1}\in{\mathfrak{S}}_{n+1}$ and $\sigma\in\operatorname{Sym}(n)$. Moreover, recall the definition of $\lambda(w,\sigma)$ given in (1.1). We start with the following two lemmas, for which we set $\overleftarrow{w}=w_{n+1}\cdots w_{1}\in{\mathfrak{S}}_{n+1}$ sending $i$ to $w_{n+2-i}$ and ${\sigma_{\uparrow}\in\operatorname{Sym}(n)}$ sending $i$ to $n+1-\sigma(i)$.
Lemma 2.4.
Let $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$. Then
$\lambda(\overleftarrow{w},\sigma_{\uparrow})=(-{\lambda_{1}},\dots,-{\lambda_{n}})\,.$
Proof.
The $i$-th cover relation in $(w,\sigma)$ merges the exact same blocks as the $i$-th cover relation in $(\overleftarrow{w},\sigma_{\uparrow})$. The merging blocks only come in opposite order, swapping the sign of $\lambda_{i}$. ∎
Lemma 2.5.
For $\lambda(w,\sigma)=(\lambda_{1},\dots,\lambda_{n})$, we have
$\big{\{}|\lambda_{1}|,\dots,|\lambda_{n}|\big{\}}=\big{\{}2,\dots,n+1\big{\}}\,.$
Proof.
This follows immediately from the observation that every number in $\{2,\dots,n+1\}$ is exactly once the maximum of the minima of two merging blocks. (One may observe in addition that the number $1$ is never such a maximum.) ∎
Combining these two lemmas yields the following symmetry on ¸$\mathcal{N}_{\mathcal{B}_{n}}$. More general forms of this symmetry have already been observed in  [MaglioneVoll, Corollary 3.3] and [poincareextended, Corollary 2.16].
Corollary 2.6.
It holds that
$\operatorname{Ino}(\overleftarrow{w},\sigma_{\uparrow})$ =$=\{2,\dots,n+1\}\setminus\operatorname{Ino}(w,\sigma)$
$\operatorname{Asc}\big{(}\lambda(\overleftarrow{w},\sigma_{\uparrow})\big{)}$ =$=\{2,\dots,n+1\}\setminus\operatorname{Asc}\big{(}\lambda(w,\sigma)\big{)}\,.$
For an entry $w_{i}\in\{1,\dots,n+1\}$ at position $i$ in the one-line notation of $w$, define two positions $\ell<i<r$ by
(2.3) $\ell=\max\left\{j<i\;\big{|}\;w_{j}<w_{i}\right\}\,,\quad r=\min\left\{j>i\;\big{|}\;w_{j}<w_{i}\right\}\,.$
In words, $\ell$ is the rightmost position left of the position $i$ such that $w_{\ell}$ is smaller than $w_{i}$. Similarly, $r$ is the leftmost position right of the position $i$ such that $w_{r}$ is smaller than $w_{i}$.
$w=\cdots\ w_{\ell}\underbrace{\phantom{sp}\cdots\phantom{sp}}_{>w_{i}}\ w_{i}\ \underbrace{\phantom{sp}\cdots\phantom{sp}}_{>w_{i}}w_{r}\ \cdots$
Observe that $\ell=-\infty$ if and only if $w_{i}$ is a left-to-right minimum of $w$, and $r=\infty$ if and only if $w_{i}$ is a right-to-left minimum.
We also record the cover relation ranks in $\Sigma_{n}$ in which the blocks containing $w_{\ell}$ and $w_{i}$, respectively the blocks containing $w_{r}$ and $w_{i}$ are merged. In symbols,
$\ell_{\sigma}$ $=\max\left\{k\in\{1,\dots,n\}\;\big{|}\;\sigma_{k}\in\{\ell,\dots,i-1\}\right\}\,,$
$r_{\sigma}$ $=\max\left\{k\in\{1,\dots,n\}\;\big{|}\;\sigma_{k}\in\{i,\dots,r-1\}\right\}\,.$
We call $\ell_{\sigma}$ the last left bar position of $w_{i}$ and $r_{\sigma}$ the last right bar position of $w_{i}$. If either of the two sets in (2.3) is empty, the corresponding last (left or right) bar position is set to be infinity. The reason for this choice of convention is that $\ell_{\sigma}$ is the rank in which $w_{i}$ is merged with something smaller in $w$ to the left, and thus set to be infinity if $w_{i}$ is a left-to-right minimum. The same reason holds for $r_{\sigma}$ if $w_{i}$ is a right-to-left minimum.
Example 2.7.
Let $w=215463\in{\mathfrak{S}}_{6}$ and $\sigma=14253\in\operatorname{Sym}_{5}$. The corresponding maximal chain in $\Sigma_{n}$ and its edge-labeling are
$2|1|5|4|6|3$ $\prec$ $12|5|4|6|3$ $\prec$ $12|5|46|3$ $\prec$ $125|46|3$ $\prec$ $125|346$ $\prec$ $123456\,,$
$\lambda(w,\sigma)=($$2$ $\overline{2},$ $6,$ $5,$ $\overline{4},$ $3$ ).))\,.) .$123456\,.$
The left-to-right minima of $w$ are $\{2,1\}$ and the right-to-left minima are $\{3,1\}$.
In the picture below, we label the bars in $w$ by $\sigma^{-1}$, meaning that the bars are removed in this order. For each entry $w_{i}$, the last left bar position $\ell_{\sigma}$ and the last right bar position $r_{\sigma}$ are written above in red, with $\ell_{\sigma}$ on the left and $r_{\sigma}$ on the right.
The following lemma follows immediately from the above observations.
Lemma 2.8.
Let $\lambda(w,\sigma)=(\lambda_{1},\dots\lambda_{n})$, let $j\in\{1,\dots,n\}$ and $i\in\{1,\dots,n+1\}$. Moreover, set $\ell_{\sigma}$ and $r_{\sigma}$ to be the last (left and, respectively, right) bar positions of $w$ for the entry $w_{i}$. Then
$\lambda_{j}=\begin{cases}w_{i}&\text{if }j=\ell_{\sigma}<r_{\sigma}\,,\\[5.0pt] \overline{w_{i}}&\text{if }j=r_{\sigma}<\ell_{\sigma}\,.\end{cases}$
Example 2.9.
We continue the previous example. For each entry $w_{i}$, we write above in red the last left and right bar positions. We moreover circle their respective minima.
Using Lemma 2.8, we recompute
$\lambda(215463,14253)=(\lambda_{1},\dots,\lambda_{5})=(\overline{2},6,5,\overline{4},3)\,.$
In particular, every position in $\{1,\dots,5\}$ is circled exactly once in agreement with Lemma 2.5, and the signs of $\lambda$ are determined whether the last (left or right) bar position is smaller.
2.4. The poset $\boldsymbol{P_{w,Y}}$
Throughout this section, fix $w\in{\mathfrak{S}}_{n+1}$ and $Y\subseteq\{1,\dots,n\}$ such that
  • $Y$ does not contain any left-to-right minima of $w$, while
  • $Y$ contains all right-to-left minima of $w$, except the number $1$.
Based on the properties of the edge-labeling $\lambda(w,\sigma)$ provided in the previous section, we define a poset $P=P_{w,Y}$ on the set $\{1,\dots,n\}$ and an injective vertex-labeling $\Lambda$ of $P$. This poset is built so that
$\mathsf{Lin}(P)=\{\sigma\in\operatorname{Sym}(n)\mid\operatorname{Ino}(w,\sigma)=Y\}\,,$
where $\mathsf{Lin}(P)$ denotes the set of all linear extensions. In particular, this implies that $\operatorname{Sym}(n)$ is equal to the disjoint union of the linear extensions of the posets $P_{w,Y}$ for all possible subsets $Y\subseteq\leavevmode\nobreak\ \{1,\dots,n\}$ with the two given properties.
We construct the poset $P$ together with its injective vertex-labeling $\Lambda$ by iterating through the numbers $n+1,\dots,2$ in this order. At each iteration step $w_{i}\in\{n+1,\dots,2\}$, we set the label $\Lambda$ of one vertex. If $w_{i}$ is neither a left-to-right nor right-to-left minimum of $w$, we add an additional cover relation to $P$, based on the already added cover relations.
  1. 1.
    If $w_{i}$ is not a left-to-right minimum, we have $\ell\geq 1$ and the set $\{\ell,\dots,i-1\}$ forms a connected subposet of $P$ with unique maximum. Denote this maximum by $a$.
  2. 2.
    If $w_{i}$ is not a right-to-left minimum, we have $r\leq n$ and the set $\{i,\dots,r-1\}$ forms a connected subposet of $P$ with unique maximum. Denote this maximum by $b$.
  3. 3.
    If $w_{i}$ is a left-to-right minimum, it is not a right-to-left minimum. Set $\Lambda(b)=\overline{w_{i}}$.
  4. 4.
    If $w_{i}$ is a right-to-left minimum, it is not a left-to-right minimum. Set $\Lambda(a)=w_{i}$.
  5. 5.
    If $w_{i}$ is neither a left-to-right nor a right-to-left minimum, add the cover relation together with the vertex label of the minimum given by
[…]
Observe that this iterative process is well-defined since the constructed Hasse diagram is a forest at every step.
Example 2.10.
We continue the previous example with $w=215463$, and let $Y=\{3,5,6\}$ satisfying the two required properties
$\{3\}\subseteq Y\subseteq\{3,4,5,6\}\,.$
The iterative construction of the poset $P$ is indicated in the following figure row-by-row and in each row left-to-right. At each step, there is set one additional vertex-labeling $\Lambda$, and in the first three iteration steps, there is also a cover relation added. This corresponds to the observation that the entries $6,5,4$ in $w$ are neither left-to-right nor right-to-left minima, while $3,2$ are.
Proposition 2.11.
Let $\sigma\in\operatorname{Sym}(n)$ and let $Y=\operatorname{Ino}(w,\sigma)$. Then, $\sigma\in\mathsf{Lin}(P_{w,Y}).$ Moreover,
$\lambda(w,\sigma)=\Lambda(\sigma)=\Big{(}\Lambda\big{(}\sigma(1)\big{)},\dots,\Lambda\big{(}\sigma(n)\big{)}\Big{)}\,,$
where $\Lambda$ is the injective vertex-labeling of $P_{w,Y}$.
Proof.
If an iteration step $w_{i}\in\{n+1,\dots,2\}$ introduces a new cover relation, the two indices $a$ and $b$ are given by $a=\sigma(\ell_{\sigma})$ and $b=\sigma(r_{\sigma})$. Then, $a\prec_{P}b$ if $\ell_{\sigma}<r_{\sigma}$ and $b\prec_{P}a$ if $r_{\sigma}<\ell_{\sigma}$. In particular, $a\prec_{P}b$ if and only if $\sigma^{-1}(a)<\sigma^{-1}(b)$. We conclude that $\sigma\in\mathsf{Lin}(P)$. The second property follows with Lemma 2.8. ∎
Example 2.12.
In our running example, $\sigma=14253\in\mathsf{Lin}(P)$ is a linear extension of $P$ and we obtain
$\Lambda(\sigma)=(\overline{2},6,5,\overline{4},3)=\lambda(w,\sigma)\,.$
2.5. Quasisymmetric functions and a proof of the main result
In order to prove Theorem 2.2, we use the following implication of the theory of reverse $(P,\omega)$-partitions from [MR4621625, §7.19].
Proposition 2.13.
Let $P$ be a poset on $\{1,\dots,n\}$ and let $\omega,\omega^{\prime}:P\to\mathbb{Z}$ be two injective vertex-labelings having the same relative orders on the cover relations. This is, for each cover relation $a\prec_{P}\leavevmode\nobreak\ b$ we have
$\omega(a)<\omega(b)\Leftrightarrow\omega^{\prime}(a)<\omega^{\prime}(b)\,.$
Then, $\omega$ and $\omega^{\prime}$ have the same descent set generating function on the set of linear extensions,
$\sum_{\sigma\in\mathsf{Lin}(P)}\boldsymbol{t}^{\operatorname{Des}(\omega(\sigma))}=\sum_{\sigma\in\mathsf{Lin}(P)}\boldsymbol{t}^{\operatorname{Des}(\omega^{\prime}(\sigma))}\,,$
where $\operatorname{Des}\big{(}\omega(\sigma)\big{)}$ denotes the descent set of $\big{(}\omega(\sigma(1)),\dots,\omega(\sigma(n))\big{)}$ and analogously for $\operatorname{Des}(\omega^{\prime}(\sigma))$.
Before proving this proposition, we deduce Theorem 2.2 as follows.
Lemma 2.14.
For a cover relation $a\prec_{P}b$ in $P=P_{w,Y}$, we have
$a<b\Longleftrightarrow\Lambda(a)>\Lambda(b)\,,$
where both comparisons are as integers.
Proof.
Within the construction of $P$, a cover relation $a\prec_{P}b$ is introduced for $a<b$ if $a$ and $b$ are the last left and right bars for some $w_{i}\in\{n+1,\dots,2\}$ and $\Lambda(a)=w_{i}$. If a cover relation $a\prec_{P}b$ is introduced for $b<a$ instead, the label of $a$ is set to $\Lambda(a)=\overline{w_{i}}$. In both cases, $\Lambda(b)$ will be determined at a later iteration step. Thus, we have $\overline{w_{i}}<\Lambda(b)<w_{i}$, so that $\Lambda(a)>\Lambda(b)$ if $a<b$, and $\Lambda(a)<\Lambda(b)$ if $a>b$ as desired. ∎
Proof of Theorem 2.2.
Using Proposition 2.11 and Lemma 2.14, it follows with Proposition 2.13 that
(2.4) $\sum_{\sigma\in\mathsf{Lin}(P_{w,Y})}\boldsymbol{t}^{\operatorname{Des}(\sigma)}=\sum_{\sigma\in\mathsf{Lin}(P_{w,Y})}\boldsymbol{t}^{\operatorname{Asc}(\lambda(w,\sigma))}\,.$
Since $\operatorname{Sym}(n)$ is the disjoint union over all linear extensions of the posets $P_{w,Y}$, the observation that $\operatorname{Ino}(w,\sigma)=Y$ for $\sigma\in\mathsf{Lin}(P_{w,Y})$ applied to (2.4) implies
$\sum_{\sigma\in\operatorname{Sym}(n)}\boldsymbol{y}^{\operatorname{Ino}(w,\sigma)}\cdot\boldsymbol{t}^{\operatorname{Asc}(\lambda(w,\sigma))}$ $=\sum_{Y}\boldsymbol{y}^{Y}\sum_{\sigma\in\mathsf{Lin}(P_{w,Y})}\boldsymbol{t}^{\operatorname{Asc}(\lambda(w,\sigma))}$
$=\sum_{Y}\boldsymbol{y}^{Y}\sum_{\sigma\in\mathsf{Lin}(P_{w,Y})}\boldsymbol{t}^{\operatorname{Des}(\sigma)}=\sum_{\sigma\in\operatorname{Sym}(n)}\boldsymbol{y}^{\operatorname{Ino}(w,\sigma)}\cdot\boldsymbol{t}^{\operatorname{Des}(\sigma)}\,,$
where the two outer sums on the right-hand side range over all subsets $Y\subseteq\{1,\dots,n\}$ that do not contain left-to-right minima of $w$ while containing all right-to-left minima of $w$ except the number $1$. ∎
We are left to provide a proof of Proposition 2.13. Let $P$ be a finite poset on $\{1,\dots,n\}$ and let $\omega:P\to\mathbb{Z}$ be an injective vertex-labeling of $P$. Following [MR4621625, §7.19], a reverse $\boldsymbol{(P,\omega)}$-partition is a map $\sigma:P\to\mathbb{N}$ satisfying
  1. 1.
    if $a\prec_{P}b$, then $\sigma(a)\leq\sigma(b)$, and
  2. 2.
    if $a\prec_{P}b$ and $\omega(a)>\omega(b)$, then $\sigma(a)<\sigma(b)$.
Denote the set of reverse $(P,\omega)$-partitions by $\mathcal{A}^{r}(P,\omega)$, and define the quasisymmetric function
$K_{P,\omega}=\sum_{\sigma\in\mathcal{A}^{r}(P,\omega)}\boldsymbol{x}^{\sigma}\in\mathbb{N}[\![\boldsymbol{x}]\!]=\mathbb{N}[\![x_{1},x_{2},x_{3},\dots]\!]\,,$
where $\boldsymbol{x}^{\sigma}=\prod_{a\in P}x_{\sigma(a)}$. Observe that $K_{P,\omega}$ only depends on the relative order of $\omega$ on the cover relations of $P$. We exploit this property to prove Proposition 2.13.
There is an alternative formula for $K_{P,\omega}$ in terms of linear extensions and the fundamental quasisymmetric functions
$L_{\operatorname{Des}(\sigma)}=\sum x_{i_{1}}\cdots\ x_{i_{n}}\,,$
where the sum ranges over all sequences $1\leq i_{1}\leq\cdots\leq i_{n}\leq n$ with $i_{k}<i_{k+1}$ if $k\in\operatorname{Des}(\sigma)$. The quasisymmetric function $K_{P,\omega}$ then has the description
(2.5) $K_{P,\omega}=\sum_{\sigma\in\mathsf{Lin}(P)}L_{\operatorname{Des}\omega(\sigma)}\,,$
see [MR4621625, Corollary 7.19.5].
Proof of Proposition 2.13.
Let $\omega,\omega^{\prime}:P\rightarrow\mathbb{Z}$ be two injective vertex-labelings such that for each cover relation $a\prec_{P}b$, we have
$\omega(a)<\omega(b)\Longleftrightarrow\omega^{\prime}(a)<\omega^{\prime}(b)\,.$
Since $K_{P,\omega}$ only depends on the relative orders of $\omega$ on the cover relations, we obtain $K_{P,\omega}=K_{P,\omega^{\prime}}$. It thus follows from (2.5) that
$\sum_{\sigma\in\mathsf{Lin}(P)}L_{\operatorname{Des}\omega(\sigma)}=\sum_{\sigma\in\mathsf{Lin}(P)}L_{\operatorname{Des}\omega^{\prime}(\sigma)}\,.$
Since $L_{\operatorname{Des}(\sigma)}$ only depends on the descent set of $\sigma$, this implies Proposition 2.13. More precisely, for any subsets $A,B\subseteq\operatorname{Sym}(n)$, the equalities of fundamental quasisymmetric function generating functions and of descent set generating functions are equivalent,
$\sum_{\sigma\in A}L_{\operatorname{Des}(\sigma)}=\sum_{\sigma\in B}L_{\operatorname{Des}(\sigma)}\quad\Longleftrightarrow\quad\sum_{\sigma\in A}\boldsymbol{t}^{\operatorname{Des}(\sigma)}=\sum_{\sigma\in B}\boldsymbol{t}^{\operatorname{Des}(\sigma)}\,.\qed$
[…]